SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Thursday, March 20, 2014

Status Report From: NOAA Office of Satellite Operations
Posted: Thursday, March 20, 2014

Customers or users who want to call to ask about problems or report outages can use the following contacts:

– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-683-1400

– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 079

GOES 
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health &
safety. For more GOES information please see:
http://www.ospo.noaa.gov/Operations/GOES/index.html 


POES 
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more POES information please see:
http://www.ospo.noaa.gov/Operations/POES/index.html 


DMSP 
OPS47 rev#98026 12:51z Marginal: Took 4 attempts to load EPHEM; was finally loaded along with stars, but missed update time. REEF added up for engineers.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety. 
For more DMSP Information please see: 
http://ngdc.noaa.gov/eog/dmsp.html


JASON-2 
All Jason-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.


SNPP
Note: SNPP rev#12391 1819z Failed: Some TLM not received and CMDING failed due to VCID “0” not being received. AFWA and NESDIS IDPS nominal, GSE will recycle MMC in back orbit. This is a known on-going issue.

All other SNPP operations were nominal over the past 24 hours. 
No anomalies affecting product delivery or s/c health & safety.
For more SNPP Information please see:
http://npp.gsfc.nasa.gov/ 


ESPC reported data outages or delays 
Fairbanks CDA: 0145Z -- UPDATE ON THE FOLLOWING: FCDAS 13M B/2 is YELLOW for right hand (RH) X-Band receive. The RH signal is degraded by approximately 20bd. Left X-BAND receive is functional. Troubleshooting has revealed the problem to likely be in the x-band waveguide portion of the receive path. New x-band waveguide switch on priority 1 order as of COB today. Test-inject circuit repaired and now functioning similar to other 13Meter systems.

 

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter